You are here: Foswiki>Commissioning Web>Com20110704 (04 Jul 2011, AndrewRakich)Edit Attach

Thermal cycling primarys with LBC FPIA

Observer(s): A. Rakich, K. Boutsia and O. Kuhn at start of the night (LBTO)
Telescope Operator: S. Allanson (LBTO)
Telescope Support: M. Midkiff (LBTO)
Instrument Support: none


Thermal cycling primary mirrors while running FPIA all night.


03:45 Had to load global Z offsets -1.8 mm Z (SX) and -1 Z (DX) before running FPIA. Then FPIA converged in several iterations on both sides. Running RangeBal.


All times are UTC unless otherwise noted.

04:10 Driving primary mirrors down in temperature and running FPIA with 10 nm convergence criterion. Fitting ~ 12 pupils in this field. Settling in for the long haul.

412 Blue pointing reference pupil centres:

a) (1771, 212), (1676, 949), (332, 1186).

4:20 LBC Red controller error, restarting cameras.

6:47 C00 REd

06:51 C00 blue.

In both cases we got this sudden large build up of Z11/Z22, don't know why. After clearing forces looked good again.

10:00 started warming the mirrors after the last entry, and haven't had a repeat of that funny Z11/Z22 build up.

10:20 closing for humidity

10:40 humidity 99%

11:15 humidity 85 and cloud clearing, reopening for some closing data points.

11:25 dawn, closing.

-- AndrewRakich - 04 Jul 2011
Topic revision: r3 - 04 Jul 2011, AndrewRakich
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback